# Exponential Runge–Kutta methods of collocation type for parabolic equations with time-dependent delay

### v1

@@ 3 ERKC methods
$V_{n+1}$ $=\mathrm{e}^{-h_{n+1}A}V_{n}\!\!\!\!\!\!\!\!\!\!$ $\!\!\!\!\!\!\!\!+h_{n+1}b_{1}(1;-h_{n+1}A)g(t_{n,1},V_{n,1},V^{\tau}(t_{n,1}))$
$\!\!\!\!\!\!\!\!+h_{n+1}b_{2}(1;-h_{n+1}A)g(t_{n+1},V_{n+1},V^{\tau}(t_{n+1})),$
$V_{n,1}$ $=\mathrm{e}^{-\tfrac{1}{3}h_{n+1}A}V_{n}\!\!\!\!\!\!\!\!\!\!$ $\!\!\!\!\!+h_{n+1}b_{1}(\tfrac{1}{3};-h_{n+1}A)g(t_{n,1},V_{n,1},V^{\tau}(t_{n,1}))$
$\!\!\!\!\!+h_{n+1}b_{2}(\tfrac{1}{3};-h_{n+1}A)g(t_{n+1},V_{n+1},V^{\tau}(t_{n+1})),$
where $V(t)=\mathrm{e}^{-\theta h_{k+1}A}V_{k}+h_{k+1}b_{1}(\theta;-h_{k+1}A)G_{k,1}^{\rm C}+h_{k+1}b_{2}(\theta;-h_{k+1}A)G_{k,2}^{\rm C},$ for $t=t_{k}+\theta h_{k+1}\in I_{k+1}$, $10\leq k\leq n-$ 1.
@@ 4 Error analysis
$U(t_{n}+\theta h_{n+1})$ $=\mathrm{e}^{-\theta h_{n+1}A}U_{n}+h_{n+1}\sum_{i=1}^{s}b_{i}(\theta;-h_{n+1}A)G^{\mathrm{I}}_{n,i},\quad 0\leq\theta\leq 1.$ (4.15)
@@
Henceforth, we suppose that (1.1) admits a solution $u:[0,T]\to V$ that is sufficiently smooth in each segment $\mathcal{I}^{\mu}$, with its derivatives belonging to $V$. Assuming that the nonlinearity $g(t,v,w)$ is sufficiently often Fréchet differentiable in a strip along the exact solution, that the delay $\tau(t)$ is sufficiently smooth in $[0,T]$ and that $\phi(t)$ is sufficiently smooth in $[-\tau_{0},0]$, it follows that the composition
@@
respectively. Subtracting (4.19) from (4.18) and noting that the order conditions
@@
hold (a consequence of (3.11) and (3.14)), we have
$\delta_{n+1}(\theta)=$ $\int_{0}^{\theta h_{n+1}}\mathrm{e}^{-(\theta h_{n+1}-\sigma)A}\int_{0}^{\sigma}\frac{(\sigma-\xi)^{s-1}}{(s-1)!}f^{(s)}(t_{n}+\xi)~{}\mathrm{d}\xi~{}\mathrm{d}\sigma$
@@
An estimate of $\delta_{n+1}(\theta)$ is given by
$h_{n+1}^{\alpha}\max_{0\leq\theta\leq 1}\|\delta_{n+1}(\theta)\|_{V}+\max_{0\leq\theta\leq 1}\|{A}^{-\alpha}\delta_{n+1}(\theta)\|_{V}\leq Ch_{n+1}^{s}\int_{t_{n}}^{t_{n+1}}\|f^{(s)}(\sigma)\|_{X}~{}\mathrm{d}\sigma.$ (4.22)
Let $e(t)=U(t)-u(t)$ denote the difference between the auxiliary function (4.15) and the exact solution. Subtracting (4.17) from (4.15) gives the error recursion
@@
$\Pi_{h}u(t_{k})=u(t_{k}),\quad\Pi_{h}u(t_{k+1})=u(t_{k+1}),\quad\mbox{and}\quad\Pi_{h}u(t_{k,i})=u(t_{k,i})~{}~{}\mbox{for}~{}i=1,2,\ldots,s.$ (4.24)
@@
provided that $u^{(s)}\in L^{\infty}((t_{k},t_{k+1});V)$, where the constant $C$ is independent of $I_{k+1}$ and the step size sequence.
@@
$\max_{0\leq t\leq T}\|\Pi_{h}U(t)-u(t)\|_{V}\leq Ch^{s}\bigg{(}\sup_{\sigma\in[0,T]}\|f^{(s)}(\sigma)\|_{X}+\sup_{\sigma\in[0,T]}\|u^{(s)}(\sigma)\|_{V}\bigg{)},$ (4.27)
@@
The proof proceeds by induction on the intervals $[0,\xi_{\mu}]$, $\mu=1,2,\ldots,m+1$. We first consider the interval $[0,\xi_{1}]$ in which case the problem reduces to a parabolic problem without delay. When the mesh point $t_{n}$ belongs to $[0,\xi_{1})$, we arrive at
@@
$\quad+C\sum_{j=1}^{n-1}(t_{n}-t_{j})^{-\alpha}h_{j}\|A^{-\alpha}\delta_{j}(1)/h_{j}\|_{V}+\|\delta_{n}(1)\|_{V}.$
Using (4.22), (4.30) and the fact that $\Pi_{h}U(t)=u(t)$ for $t\in[-\tau_{0},0]$, we get by applying the discrete Gronwall lemma (Lemma 4.1) the following result: for any mesh point $t_{n}$ belonging to $[0,\xi_{1}]$, it holds
@@
For a mesh point $t_{n}$ belonging to $[\xi_{\mu},\xi_{\mu+1})$, from (4.29) and (4.35), we obtain
@@
$\quad-f(t_{j-1,i})\Big{)}-\delta_{j}(1)\bigg{)},$
where $t_{\chi(\xi_{\mu})}=\xi_{\mu}$. Now, estimating this term in $V$ similarly to (4.31), using (4.35) and (4.36) and applying the discrete Gronwall lemma (Lemma 4.1) leads to the desired result: for a mesh point $t_{n}$ belonging to $[\xi_{\mu},\xi_{\mu+1}]$,
@@
The existence and uniqueness of the numerical solutions $U_{n,i}~{}(i=1,2,\ldots,s)$ can be verified by a standard induction argument and fixed-point iteration. Note that the Lipschitz constant $L$ of $g$ can be controlled for sufficiently small $h$ (see (4.28) and (4.29)). ∎
@@
${h_{j}}\leq\varrho{h_{j+1}},\quad a_{n,j}\leq a(t_{n}-t_{j})^{\gamma}\quad\mbox{and}\quad b_{j}\leq bt_{j}^{\delta},$
where $\varrho>1$, $a>0$, $b>0$, $-1<\gamma\leq 0$ and $-1<\delta\leq 0$. Further assume that the nonegative sequence $\{\varepsilon_{j}\}_{j=1}^{N}$ satisfies the inequalities
$\varepsilon_{n}\leq\sum_{j=1}^{n-1}h_{j}a_{n,j}\varepsilon_{j}+b_{n},\quad\mbox{for}~{}1\leq n\leq N.$
@@
$\varepsilon_{n}\leq Cbt_{n}^{\delta},\quad\mbox{for}~{}1\leq n\leq N,$
where the constant $C$ depends on $\varrho$, $a$, $\gamma$, $\delta$ and on $t_{N}$.
The convergence result for ERKC-C methods is presented in the next theorem. The construction of $U(t)$ (4.15) and $V(t)$ (3.10) follows a similar approach with the only difference being in $G_{n,i}^{\rm I}$ and $G^{\rm C}_{n,i}$. Specifically, $G^{\rm I}_{n,i}$ incorporates $U(t)$ and its interpolation $\Pi_{h}U(t)$, while $G^{\rm C}_{n,i}$ is based on $V(t)$ only. The difference in the proof lies in the approximation of the delayed solution: for ERKC-I methods, the error in the delayed solution is estimated with the help of (4.25), whereas for ERKC-C methods, the error in the delayed solution is estimated directly through $\|V(t)-u(t)\|_{V}$. The proof closely resembles that of ERKC-I methods and is therefore skipped.
@@
Let the initial value problem (1.1) satisfy Assumptions 2.1-2.2, and consider for its numerical solution the ERKC-C method (3.12). Let the step size sequence $\{h_{j}\}_{j=1}^{N}$ satisfy the condition (4.26). If $f:\mathcal{I}^{\mu}\to X$ (4.16) is $s$ times differentiable for $\mu=1,2,\ldots,m+1$ with $f^{(s)}\in L^{\infty}(0,T;X)$, then for $h=\max_{1\leq j\leq N}h_{j}$ sufficiently small, the numerical solution (3.10) is determined uniquely and the error of $V(t)$ satisfies
$\max_{0\leq t\leq T}\|V(t)-u(t)\|_{V}\leq Ch^{s}\sup_{\sigma\in[0,T]}\|f^{(s)}(\sigma)\|_{X},$
@@ 5 Superconvergence
and that $f(t)$ has additional smoothness (see Theorem 5.1). Condition (5.39) is satisfied, for example, when the collocation parameters are chosen as Radau points with $s\geq 2$ or as Gauss points. We shall extend this superconvergence result to ERKC-I methods for semilinear parabolic problems with time-dependent delay (1.1). The following improved error estimate of $\Pi_{h}U(t)$ in each interval $I_{k+1}$ is required (cf. (4.25)),
@@
$\quad+Ch^{s+1}\sup_{\sigma\in[0,T]}\|u^{(s+1)}(\sigma)\|_{V},$
where $C_{\mathrm{S}}$ is given by Lemma 5.1 below. The constant $C$ depends on $T$, but is independent of the step size sequence.
@@
$\delta_{n+1}(1)=h_{n+1}^{s+1}\psi_{s+1}(-h_{n+1}A)f^{(s)}(t_{n}^{+})+\tilde{\delta}_{n+1}(1),$ (5.41)
@@
$\Bigg{\|}\sum_{j=1}^{n}\mathrm{e}^{-(t_{n}-t_{j})A}\tilde{\delta}_{j}\Bigg{\|}_{V}\leq C\sum_{j=1}^{n}\|\hat{\delta}_{j}\|_{V}\leq Ch^{s+1}\int_{0}^{t_{n}}\|f^{(s+1)}(\sigma)\|_{V}~{}\mathrm{d}\sigma.$ (5.42)
Using Lemma 5.1, we further get
@@
Combining (5.42) with (5.43), we obtain
@@
Remark 5.1.
In particular, if $X$ is a separable Hilbert space and $A$ is linear, self-adjoint, positive definite with a compact inverse (defined in ${D}(A)\subset X$), then (5.43) is satisfied with $C_{\mathrm{S}}$ independent of $n$ and the step size sequence.
@@
where $C_{\mathrm{S}}\leq C_{0}\ln\frac{T}{\min_{1\leq j\leq N}h_{j}}$ and the constants $C$ and $C_{0}$ are independent of $n$ and the step size sequence.
@@
A combination of (5.46)-(5.48) leads to
$\|S(n)\|_{V}\leq CC_{\mathrm{S}}h^{s+1}\cdot\max_{0\leq j\leq n-1}\|w_{j}\|_{V},\quad\mbox{with}~{}C_{\mathrm{S}}\leq C_{0}\frac{T}{\min_{1\leq j\leq N}h_{j}},$ (5.49)
where the constant $C_{0}$ is independent of $n$ and the step size sequence. ∎
@@
In many situations, the number $C_{\mathrm{S}}\leq C_{0}\ln\frac{T}{\min_{1\leq j\leq N}h_{j}}$ in (5.49) can be controlled independently of $\min_{1\leq j\leq N}h_{j}$. For example, consider a constrained mesh $I_{h}=\{nh:n=0,1,\ldots,N\}\cup\{\xi_{1},\xi_{2},\ldots,\xi_{m}\}$, which is generated by including the potential discontinuity points into the uniform partition. Similar to the derivation of (5.49), it follows that $\|S(N)\|_{V}\leq Ch^{s+1}(\ln N+1)\max_{0\leq j\leq N-1}\|w_{j}\|_{V}$.
@@
Let the initial value problem (1.1) satisfy Assumptions 2.1-2.2, and consider for its numerical solution the ERKC-C method (3.12) whose underlying quadrature rule is of order $s+1$. Let the step size sequence $\{h_{j}\}_{j=1}^{N}$ satisfy the condition (4.26). If $f:\mathcal{I}^{\mu}\to V$ is $s+1$ times differentiable for $\mu=1,2,\ldots,m+1$ with $f^{(s)}\in L^{\infty}(0,T;V)$, $f^{(s+1)}\in L^{1}(0,T;V)$, then for $h=\max_{1\leq j\leq N}h_{j}$ sufficiently small the error of $V(t)$ satisfies
$\max_{t\in[0,T]}\|V(t)-u(t)\|_{V}$ $\leq CC_{\mathrm{S}}h^{s+1}\bigg{(}\int_{0}^{T}\|f^{(s+1)}(\sigma)\|_{V}\mathrm{~{}d}\sigma+\sup_{\sigma\in[0,T]}\|f^{(s)}(\sigma)\|_{V}\bigg{)},$
Theorems 5.1 and 5.2 give superconvergence results of order $s+1$. From Theorem 6 in [11], the convergence order of ERKC methods with constant step size further achieves $s+1+\beta$ at the nodes within $[0,\xi_{1}]$, provided that the underlying quadrature rule is of order $s+2$ and some slightly stronger assumptions regarding spatial regularity hold. The extension of this result to ERKC-I and ERKC-C methods over $[0,T]$ is not feasible under the present framework. To address this limitation, we propose the following modified ERKC-I method by adjusting the interpolation approach,
@@
$U_{n,i}$ $=\mathrm{e}^{-c_{i}h_{n+1}A}U_{n}+h_{n+1}\sum_{j=1}^{s}a_{ij}(-h_{n+1}A)G^{\rm I}_{n,j},\quad i=1,2,\ldots,s,$
$G^{\rm I}_{n,i}$ $=g(t_{n,i},U_{n,i},(\Pi_{h}^{*}U)^{\tau}(t_{n,i})),\quad i=1,2,\ldots,s.$
For $t\in[t_{n},t_{n+1}]$, we have that $\hat{t}=t-\tau(t)$ lies either in $[-\tau_{0},0]$ or in $\mathcal{I}^{\mu}$ for some $0\leq\mu\leq m$. We set $\Pi^{*}_{h}U(\hat{t})=\phi(\hat{t})$ for $\hat{t}\in[-\tau_{0},0]$. For $\hat{t}\in I_{k+1}\subseteq\mathcal{I}^{\mu}$, the value of $\Pi^{*}_{h}U(\hat{t})$ is given by a polynomial based on the interpolation at mesh points as follows:
@@
$P_{k+1}(t_{q})=U_{q}\quad\mbox{for}~{}t_{q}\in S^{*}_{k+1}=\{t_{k-\ell_{k}+1},t_{k-\ell_{k}+2},\ldots,t_{k},t_{k+1},\ldots,t_{k+r_{k}-1},t_{k+r_{k}}\}\subseteq\mathcal{I}^{\mu}.$ (5.51)
@@
The number $C$ is independent of $I_{k+1}$ and the step size sequence if the mesh $I_{h}$ is quasi-uniform. Moreover, we need the following assumptions on spatial regularity.
@@
Let $g:[0,T]\times V\times V\rightarrow X$ be Fréchet differentiable with respect to the second and the third variable with locally Lipschitz-continuous derivative. Further, let $\beta\in[0,1]$ be such that
@@
$\bigg{\|}g(t,v,w)-g(t,u(t),w)-\frac{\partial g}{\partial u}(t,u(t),u^{\tau}(t))(v-u(t))\bigg{\|}_{X}$ (5.54)
$\qquad\leq\widetilde{L}\big{(}\|v-u(t)\|_{V}+\|w-u^{\tau}(t)\|_{V}\big{)}\|v-u(t)\|_{V},$
for all $t\in[0,T]$ and all $v,w\in V$ with
$\max(\|v-u(t)\|_{V},\|w-u^{\tau}(t)\|_{V})\leq R.$

### v3

@@ 3 ERKC methods
$W_{n+1}$ $=\mathrm{e}^{-h_{n+1}A}W_{n}\!\!\!\!\!\!\!\!\!\!$ $\!\!\!\!\!\!\!\!+h_{n+1}b_{1}(1;-h_{n+1}A)g(t_{n,1},W_{n,1},W^{\tau}(t_{n,1}))$
$\!\!\!\!\!\!\!\!+h_{n+1}b_{2}(1;-h_{n+1}A)g(t_{n+1},W_{n+1},W^{\tau}(t_{n+1})),$
$W_{n,1}$ $=\mathrm{e}^{-\tfrac{1}{3}h_{n+1}A}W_{n}\!\!\!\!\!\!\!\!\!\!$ $\!\!\!\!\!+h_{n+1}b_{1}(\tfrac{1}{3};-h_{n+1}A)g(t_{n,1},W_{n,1},W^{\tau}(t_{n,1}))$
$\!\!\!\!\!+h_{n+1}b_{2}(\tfrac{1}{3};-h_{n+1}A)g(t_{n+1},W_{n+1},W^{\tau}(t_{n+1})),$
where $W(t)=\mathrm{e}^{-\theta h_{k+1}A}W_{k}+h_{k+1}b_{1}(\theta;-h_{k+1}A)G_{k,1}^{\rm C}+h_{k+1}b_{2}(\theta;-h_{k+1}A)G_{k,2}^{\rm C},$ for $t=t_{k}+\theta h_{k+1}\in I_{k+1}$, $0\leq k\leq n-1$.
@@ 4 Error analysis
$U(t_{n}+\theta h_{n+1})=\mathrm{e}^{-\theta h_{n+1}A}U_{n}+h_{n+1}\sum_{i=1}^{s}b_{i}(\theta;-h_{n+1}A)G^{\mathrm{I}}_{n,i},~~0\leq\theta\leq 1.$ (4.1)
@@
Henceforth, we suppose that (1.1) admits a solution $u:[0,T]\to V$ that is sufficiently smooth in each segment $\mathcal{I}^{\mu}$, with its derivatives belonging to $V$. Assuming that the nonlinearity $g(t,v,w)$ is sufficiently often Fréchet differentiable in a strip along the exact solution, that the delay $\tau(t)$ is sufficiently smooth in $[0,T]$ and that $\phi(t)$ is sufficiently smooth in $\tau_{0},0]$, it follows that the composition
@@
respectively. Subtracting (4.5) from (4.4) and using
@@
(which is an immediate consequence of (3.9) and (3.12)), we have
$\delta_{n+1}(\theta)=$ $\int_{0}^{\theta h_{n+1}}\mathrm{e}^{-(\theta h_{n+1}-\sigma)A}\int_{0}^{\sigma}\frac{(\sigma-\rho)^{s-1}}{(s-1)!}f^{(s)}(t_{n}+\rho)~\mathrm{d}\rho~\mathrm{d}\sigma$
@@
We note that the identities (4.6) are usually called order conditions. Using (3.13), the following estmate of $\delta_{n+1}(\theta)$ is obtained
$h_{n+1}^{\alpha}\max_{0\leq\theta\leq 1}\|\delta_{n+1}(\theta)\|_{V}+\max_{0\leq\theta\leq 1}\|\delta_{n+1}(\theta)\|_{X}\leq Ch_{n+1}^{s}\int_{t_{n}}^{t_{n+1}}\|f^{(s)}(\sigma)\|_{X}~\mathrm{d}\sigma.$ (4.8)
Let e(t)=U(t)u(t)e(t)=U(t)-u(t) denote the difference between the auxiliary function (4.1) and the exact solution. Subtracting (4.3) from (4.1) gives the error recursion
@@
$\Pi_{h}u(t_{k})=u(t_{k}),~\Pi_{h}u(t_{k+1})=u(t_{k+1}),~\mbox{and}~\Pi_{h}u(t_{k,i})=u(t_{k,i})~~\mbox{for}~1\leq i\leq s.$ (4.10)
@@
provided that $u^{(s)}\in L^{\infty}(0,T;V)$, where the constant $C$ is independent of $I_{k+1}$ and the step size sequence.
@@
$\max_{0\leq t\leq T}\|\Pi_{h}U(t)-u(t)\|_{V}\leq Ch^{s}\bigg(\sup_{\sigma\in[0,T]}\|f^{(s)}(\sigma)\|_{X}+\sup_{\sigma\in[0,T]}\|u^{(s)}(\sigma)\|_{V}\bigg),$
@@
The proof proceeds by induction on the intervals $0,\xi_{\mu}]$, $\mu=1,2,\ldots,m+1$. We first consider the interval $0,\xi_{1}]$ in which case the problem reduces to a parabolic problem without delay. When the mesh point $t_{n}$ belongs to $0,\xi_{1})$, we arrive at
@@
$\quad+C\sum_{j=1}^{n-1}(t_{n}-t_{j})^{-\alpha}\|\delta_{j}(1)\|_{X}+\|\delta_{n}(1)\|_{V}.$
Note that $h_{j}\leq\varrho h_{j+1}$ and $(t_{n}-t_{j})^{-\alpha}\leq(1+\varrho)^{\alpha}(t_{n}-t_{j-1})^{-\alpha}$. Using (4.8), (4.15) and the fact that $\Pi_{h}U(t)=u(t)$ for $t\in[-\tau_{0},0]$, we get by applying the discrete Gronwall lemma (Lemma 4.1) the following result: for any mesh point $t_{n}$ belonging to $0,\xi_{1}]$, it holds
@@
For a mesh point $t_{n}$ belonging to $\xi_{\mu},\xi_{\mu+1})$, from (4.14) and (4.20), we obtain
@@
$\qquad-f(t_{j-1,i})\Big)-\delta_{j}(1)\bigg),$
where $t_{\chi(\xi_{\mu})}=\xi_{\mu}$. Now, estimating this term in $V$ similarly to (4.16), using (4.20) and (4.21) and applying the discrete Gronwall lemma (Lemma 4.1) leads to the desired result: for a mesh point $t_{n}$ belonging to $\xi_{\mu},\xi_{\mu+1}]$,
@@
The existence and uniqueness of the numerical solutions $U_{n,i}~(n=0,1,\ldots,N-1,~i=1,2,\ldots,s)$ can be verified by fixed-point iteration (cf. [22]), proceeding by induction on the intervals $\mathcal{I}^{\mu}$. Since $g$ is locally Lipschitz continuous, the convergence proof proceeds as long as the numerical solution remains within a prescribed neighborhood of the exact solution, which can be ensured by choosing the time step size $h$ sufficiently small. ∎
@@
$a_{n,j}\leq a(t_{n}-t_{j})^{\gamma}\quad\mbox{and}\quad b_{j}\leq b,$
where $a>0$, $b>0$ and $1<\gamma\leq 0$. Further assume that the nonegative sequence $\{\varepsilon_{j}\}_{j=1}^{N}$ satisfies the inequalities
$\varepsilon_{n}\leq\sum_{j=1}^{n-1}h_{j+1}a_{n,j}\varepsilon_{j}+b_{n},\quad\mbox{for}~1\leq n\leq N.$
@@
$\varepsilon_{n}\leq Cb,\quad\mbox{for}~1\leq n\leq N,$
where the constant $C$ depends on $a$, $\gamma$ and on $t_{N}$.
The convergence result for ERKC-C methods is presented in the next theorem. The construction of $U(t)$ (4.1) and $W(t)$ (3.8) follows a similar approach with the only difference being in $G_{n,i}^{\rm I}$ and $G^{\rm C}_{n,i}$. Specifically, $G^{\rm I}_{n,i}$ incorporates $U(t)$ and its interpolation $\Pi_{h}U(t)$, while $G^{\rm C}_{n,i}$ is based on $W(t)$ only. The only difference in the proof comes from the different approximation of the delayed solution: for ERKC-I methods, the error in the delayed solution is estimated with the help of (4.11), whereas for ERKC-C methods, the error in the delayed solution is estimated directly through $\|W(t)-u(t)\|_{V}$. As the proof closely resembles that of ERKC-I methods, it is skipped.
@@
Let the initial value problem (1.1) satisfy Assumptions 2.1-2.2, and consider for its numerical solution the ERKC-C method (3.10). Let the step size sequence $\{h_{j}\}_{j=1}^{N}$ satisfy the condition (4.12). If $f:\mathcal{I}^{\mu}\to X$ (4.2) is $s$ times differentiable for $\mu=1,2,\ldots,m+1$ with $f^{(s)}\in L^{\infty}(0,T;X)$, then for $h=\max_{1\leq j\leq N}h_{j}$ sufficiently small, the numerical solution (3.8) is determined uniquely and the error of $W(t)$ satisfies
$\max_{0\leq t\leq T}\|W(t)-u(t)\|_{V}\leq Ch^{s}\sup_{\sigma\in[0,T]}\|f^{(s)}(\sigma)\|_{X},$
@@ 5 Superconvergence
and that $f(t)$ has additional smoothness (see Theorem 5.1). Condition (5.1) is satisfied, for example, when the collocation parameters are chosen as Radau points with $s\geq 2$ or as Gauss points. In this section, we show that both the ERKC-I and ERKC-C methods with underlying quadrature rule of order $s+1$ actually achieve order $s+1$ over the entire time interval $[0,T]$, rather than only at the nodes.
We shall first extend this superconvergence result to ERKC-I methods for semilinear parabolic problems with time-dependent delay (1.1). The following improved error estimate of $\Pi_{h}U(t)$ in each interval $I_{k+1}$ is required (cf. (4.11)),
@@
$\quad+Ch^{s+1}\sup_{\sigma\in[0,T]}\|u^{(s+1)}(\sigma)\|_{V}.$
In general, the size of $C_{\mathrm{S}}$ depends on the chosen step size sequence. An upper bound is provided in Lemma 5.1 below. However, when the step sizes are constant or when the operator $A$ and the space $X$ satisfy certain conditions (see Remark 5.1 below), $C_{\mathrm{S}}$ is independent of the step size sequence. On the other hand, the constant $C$ depends on $T$, but not on the step size sequence.
@@
$\delta_{j}(1)=h_{j}^{s+1}\psi_{s+1}(-h_{j}A)f^{(s)}(t_{j-1}^{+})+\widetilde{\delta}_{j}(1),$ (5.3)
@@
$\Bigg\|\sum_{j=1}^{n}\mathrm{e}^{-(t_{n}-t_{j})A}\widetilde{\delta}_{j}(1)\Bigg\|_{V}\leq C\sum_{j=1}^{n}\|\widetilde{\delta}_{j}(1)\|_{V}\leq Ch^{s+1}\int_{0}^{t_{n}}\|f^{(s+1)}(\sigma)\|_{V}~\mathrm{d}\sigma.$ (5.4)
In the case of uniform time stepping (i.e., $h_{j}=h$ for all $j$), from the proof of [13, Theorem 2], we have
$\Bigg\|\sum_{j=1}^{n}h_{j}^{s+1}\mathrm{e}^{-(t_{n}-t_{j})A}{\psi}_{s+1}(-h_{j}A)f^{(s)}(t_{j-1}^{+})\Bigg\|_{V}$
$\qquad\leq C_{\mathrm{S}}h^{s+1}\bigg(\int_{0}^{T}\|f^{(s+1)}(\sigma)\|_{V}\mathrm{~d}\sigma+\sup_{\sigma\in[0,T]}\|f^{(s)}(\sigma)\|_{V}\bigg),$ (5.5)
where $C_{\mathrm{S}}$ is a constant independent of $n$ and $h$. In the case of variable step sizes satisfying (4.12), estimate (5.5) also holds, provided that additional assumptions on $A$ and the space $X$ are satisfied (see Remark 5.1). In the absence of these assumptions, the following estimate can still be derived from Lemma 5.1
@@
Combining (5.4) with (5.5)-(5.6), we obtain
@@
where
$1\leq C_{\mathrm{S}}\leq\varrho\ln\frac{T}{\min_{1\leq j\leq N}h_{j}}+2$ (5.7)
and the constant $C$ is independent of $n$ and the step size sequence.
@@
A combination of (5.11)-(5.12) leads to, for $n\geq 3$,
$\|S(n)\|_{V}\leq CC_{n}h^{s+1}\cdot\max_{0\leq j\leq n-1}\|w_{j}\|_{V},\quad\mbox{with}~C_{n}=\varrho\ln\frac{t_{n}-t_{1}}{h_{n}}+2.$
Defining $C_{\mathrm{S}}=\sup_{t_{n}\leq T}C_{n}$, it follows that, for $n=1,2,\ldots,N$,
$\|S(n)\|_{V}\leq CC_{\mathrm{S}}h^{s+1}\cdot\max_{0\leq j\leq n-1}\|w_{j}\|_{V},\quad\mbox{with}~1\leq C_{\mathrm{S}}\leq\varrho\ln\frac{T}{\min_{1\leq j\leq N}h_{j}}+2,$
which completes the proof. ∎
@@
We now discuss the dependence of $C_{\mathrm{S}}$ on the step size sequence in the case of variable time steps satisfying (4.12). If $X$ is a separable Hilbert space and $A$ is linear, self-adjoint, positive definite with a compact inverse (defined in ${D}(A)\subset X$), by using spectral decomposition and the technique developed in [13, Theorem 2], one can prove that (5.5) is satisfied with $C_{\mathrm{S}}$ independent of $n$ and the step size sequence.
In the worst case, when the step size sequence satisfies (4.12), the number $C_{\mathrm{S}}$ in estimate (5.6) follows from Lemma 5.1. In many situations, the number $C_{\mathrm{S}}$ in (5.7) can be controlled independently of $\min_{1\leq j\leq N}h_{j}$. For example, consider a constrained mesh $I_{h}=\{nh:n=0,1,\ldots,N\}\cup\{\xi_{1},\xi_{2},\ldots,\xi_{m}\}$, which is generated by including the potential discontinuity points into the uniform partition. Similar to the derivation of (5.7), it follows that $\|S(N)\|_{V}\leq Ch^{s+1}(\ln N+1)\max_{0\leq j\leq N-1}\|w_{j}\|_{V}$.
@@
Let the initial value problem (1.1) satisfy Assumptions 2.1-2.2, and consider for its numerical solution the ERKC-C method (3.10) whose underlying quadrature rule is of order $s+1$. Let the step size sequence $\{h_{j}\}_{j=1}^{N}$ satisfy the condition (4.12). If $f:\mathcal{I}^{\mu}\to V$ is $s+1$ times differentiable for $\mu=1,2,\ldots,m+1$ with $f^{(s)}\in L^{\infty}(0,T;V)$, $f^{(s+1)}\in L^{1}(0,T;V)$, then for $h=\max_{1\leq j\leq N}h_{j}$ sufficiently small the error of $W(t)$ satisfies
$\max_{t\in[0,T]}\|W(t)-u(t)\|_{V}$ $\leq CC_{\mathrm{S}}h^{s+1}\bigg(\int_{0}^{T}\|f^{(s+1)}(\sigma)\|_{V}\mathrm{~d}\sigma+\sup_{\sigma\in[0,T]}\|f^{(s)}(\sigma)\|_{V}\bigg).$
In general, the size of $C_{\mathrm{S}}$ depends on the chosen step size sequence. An upper bound is provided in Lemma 5.1. However, when the step sizes are constant or when the operator $A$ and the space $X$ satisfy certain conditions (see Remark 5.1), $C_{\mathrm{S}}$ is independent of the step size sequence. On the other hand, the constant $C$ depends on $T$, but not on the step size sequence.
Theorems 5.1 and 5.2 give superconvergence results of order $s+1$. From Theorem 6 in [13], the convergence order of ERKC methods with constant step size further achieves $s+1+\beta$ at the nodes within $0,\xi_{1}]$, provided that the underlying quadrature rule is of order $s+2$ and some slightly stronger assumptions regarding spatial regularity hold. The extension of this result to ERKC-I and ERKC-C methods over $[0,T]$ is not feasible under the present framework. To address this limitation, we propose the following modified ERKC-I method by adjusting the interpolation approach,
@@
$U_{n,i}$ $=\mathrm{e}^{-c_{i}h_{n+1}A}U_{n}+h_{n+1}\sum_{j=1}^{s}a_{ij}(-h_{n+1}A)G^{\rm I}_{n,j},\quad 1\leq i\leq s,$
$G^{\rm I}_{n,j}$ $=g(t_{n,j},U_{n,j},(\Pi_{h}^{*}U)^{\tau}(t_{n,j})),\quad 1\leq j\leq s.$
[…]
@@
$P_{k+1}(t_{q})=U_{q}\quad\mbox{for}~t_{q}\in S^{*}_{k+1}=\{t_{k-\ell_{k}+1},\ldots,t_{k},t_{k+1},\ldots,t_{k+r_{k}}\}\subseteq\mathcal{I}^{\mu}.$ (5.14)
In practice, a symmetric (or nearly symmetric, for $s$ being odd) strategy based on $s+2$ interpolation points is used for choosing $\ell_{k}$ and $r_{k}$, with the necessary adjustments to ensure that the interpolation stencil does not cross primary discontinuities.
@@
The number $C$ is independent of $I_{k+1}$ and the step size sequence provided that the following condition holds: there are constants $\kappa_{1}$ and $\kappa_{2}$ such that
$\kappa_{1}\leq h_{j+1}/h_{j}\leq\kappa_{2}\quad\mbox{for all}~j.$ (5.16)
In particular, quasi-uniform meshes satisfy this condition.
Moreover, we need the following assumptions on spatial regularity.
@@
Let $g:[0,T]\times V\times V\rightarrow X$ be Fréchet differentiable with respect to the second variable and let $\frac{\partial g}{\partial v}(t,v,w)$ be locally Lipschitz-continuous in a strip along the exact solution. Further, let $\beta\in[0,1]$ be such that
@@
$\bigg\|g(t,v,u^{\tau}(t))-g(t,u(t),u^{\tau}(t))-\frac{\partial g}{\partial u}(t,u(t),u^{\tau}(t))(v-u(t))\bigg\|_{X}\leq\widetilde{L}\|v-u(t)\|_{V}^{2},$ (5.18)
for all $t\in[0,T]$ and all $v\in V$ with $\|v-u(t)\|_{V}\leq R$.